# Concise and Organized Perception Facilitates Reasoning in Large Language Models

### v4

@@ Abstract
Exploiting large language models (LLMs) to tackle reasoning has garnered growing attention. It still remains highly challenging to achieve satisfactory results in complex logical problems, characterized by plenty of premises within the prompt and requiring multi-hop reasoning. In particular, the reasoning capabilities of LLMs are brittle to disorder and distractibility. In this work, we first examine the mechanism from the perspective of information flow and reveal that LLMs exhibit failure patterns akin to human-like cognitive biases when dealing with disordered and irrelevant content in reasoning tasks. However, in contrast to LLMs, disordered and irrelevant content does not significantly decrease human performance, as humans have a propensity to distill the most relevant information and systematically organize their thoughts, aiding them in responding to questions. Stem from that, we further propose a novel reasoning approach named Concise and Organized Perception (COP). COP carefully analyzes the given statements to identify the most pertinent information while eliminating redundancy efficiently. It then prompts the LLMs in a more organized form that adapts to the model’s inference process. By perceiving concise and organized context, the reasoning abilities of LLMs can be better elicited. Extensive experimental results on several popular logical benchmarks (ProofWriter, PrOntoQA, PrOntoQA-OOD, and FOLIO) and math benchmark (DI-GSM) show that COP significantly outperforms previous state-of-the-art methods.
@@ 1 Introduction
The field of large language models (LLMs) has witnessed significant progress in complex reasoning with the advent of Chain-of-thought (CoT) prompting [41] and a series of related works [31, 26, 45, 43, 3]. These breakthroughs have yielded remarkable achievements in various applications, including arithmetic, commonsense, symbolic reasoning, etc. [36, 1, 6, 20, 15, 5], and have sparked widespread enthusiasm within the community to continuously explore the immense potential of LLMs in tackling complex reasoning tasks. However, the performance of LLMs drastically decreases when handling intricate tasks characterized by plenty of premises within the prompt and requiring multi-hop reasoning. One primary issue is distractibility [34], where the reasoning capabilities of LLMs are highly susceptible to deterioration when confronted with irrelevant context. Another failure mode that has garnered significant attention recently is disorder [33, 7], where the performance of LLMs Severely drops when the premise order does not align with the context required in intermediate reasoning steps. Figure 1(a) presents an example from ProofWriter [35], which is a deductive logical dataset (in deductive reasoning, disorder can be further understood as that the premises are not in the same order as the ground truth proof). The complexity of the context in premises directly affects the difficulty of reasoning. When the given premises are disordered and contain much information irrelevant to the reasoning, the models face a higher risk of selecting the wrong information at some stage. This often leads to an incomplete proof and subsequently an incorrect answer. Figure 1(c) illustrates such a misleading step, where the original CoT selects the wrong reasoning path (highlighted in red). This observation indicates that LLMs usually struggle with proof planning when irrelevant and disordered content hinders, as also revealed in some contemporaneous works[2, 34, 7].
@@
Figure 1: (a) A 5-hop example of ProofWriter dataset, showcasing plenty of premises and the question to be answered. Some premises are omitted for brevity. (b) Corresponding reconstruction of concise and organized perception. Superscript serial numbers represent the logical orders according to the gold proof. The concise input contains only relevant information but lacks organizational structure. In contrast, the organized input arranges statements consistently with the gold reasoning path, albeit including some redundant information. (c) LLMs output results. (d) Results of a confirmatory experiment.
@@
In this work, we first briefly investigate the underlying mechanism of the influence of disordered and irrelevant content on reasoning. Specifically, we adopt neuron saliency score analysis, which is an important approach for pinpoint the information flow and the crucial interactions between tokens [12, 19, 40], and we conclude three phenomena. (i) LLMs struggle to identify the correct entry point of the reasoning path when faced with disorder and distraction. Intuitively, humans face the same difficulties (i.e., hard to get started) when navigating through a pile of complex information. (ii) The current step always highly focuses on the previous step, to the extent that it might even make up non existing premises to accommodate the preceding step. Consequently, it is challenging in allocating sufficient energy to identify the most accurate step. This phenomenon can be attributed to the model’s design, which employs a left-to-right reading paradigm, and closely mimics the natural process of human language comprehension. (iii) In addition, salient information flow from irrelevant information renders models prone to distractions, inadvertently causing them to focus on irrelevant content, which ultimately leads to failures in reasoning.
@@
The information flow analysis reveals how irrelevant and disorganized information can affect model reasoning, and suggests the inertia of LLMs when tackling complex tasks is very similar to the one during human problem-solving process [17, 23]. However, in contrast to LLMs, the disordered and irrelevant content does not significantly decline human performance, which benefits from the fact that humans tend to distill the most relevant information and organize their thoughts in an orderly manner, such as constructing a mind map, in advance. This allows them to address the question more more quickly and accurately by referring to the mind map [16, 14].
@@
Arise from that, we propose a novel reasoning approach named Concise and Organized Perception (COP). Specifically, COP initially performs capturing of locally-related premise segments among the given premises, with the intent to facilitate an initial comprehension of the input context. Next, COP leverages the query question as an anchor to integrate relevant pieces of locally-related premises generated by the first step, creating a tree-like mind map structure that presents global information in an orderly manner and can eliminate irrelevant information. Subsequently, LLMs are prompted by the reconstructed context, which are organized in a progressively ordered manner from the mind map to better adapt to the inference process of the model. We believe such reconstruction perceives more concise and organized information, which noticeably reduces the difficulty of model inference and better elicits the reasoning ability. Figure 1(b)(c) shows an example where LLMs are empowered to obtain the correct answer.
@@
Meanwhile, we further conducted a simple confirmatory experiment by randomly selecting 196 samples and reconstructing the context based on the provided ground-truth proofs, as shown in Figure 1(b)  [^1] Notice that this implementation is merely demonstrative, and differs from the actual method as no ground-truth can be utilized.. The results in Figure 1(d) demonstrate that combining our approach with the CoT baseline yields a relative performance improvement of over 100% (35.9% vs 71.9%) in a 5-hop setting. The results also indicate the complementarity between concise and organized perception. Besides synthetic logical reasoning, our method achieves consistent performance improvements on both real-world (FOLIO) and mathematical (DI-GSM) reasoning benchmarks. Specially, COP surpasses the CoT baseline by 9% on the FOLIO benchmark.
@@ 2 Related work
Large language models (LLMs) have demonstrated impressive few-shot learning capabilities [4, 32, 8, 29, 37]. Recent work has shown that LLMs, combined with in-context learning (ICL) and chain-of-thought (CoT) prompting, are capable of reasoning to an extent [21, 31, 28, 41, 26, 27]. However, it still remains highly challenging to achieve satisfactory results in complex logical problems.  [42] explores the logical flaws of LLMs on logical reasoning datasets from four dimensions including answer correctness, explanation correctness, explanation completeness and explanation redundancy.  [39] proposes an automatic approach to evaluate the logical reasoning abilities of LLMs based on propositional and predicate logic, which systematically identifies poor logical rules for LLMs’ reasoning. In particular, the reasoning capabilities of LLMs are susceptible to deterioration when confronted with inputs that are either arranged in a disordered manner [33, 7] or peppered with irrelevant information [34]. Specifically,  [33] investigated how reasoning ability is affected by the traversal direction of the ontology, and  [7] found that the premise order significantly affects LLMs’ reasoning performance. In another study,  [34] observed that the performance of language models tends to decrease when irrelevant context is included in the problem statement. These phenomena aligns with the human preferences for solving logical problems [16, 14, 22]. Differing from their works, we further investigate the impact of disordered and irrelevant content on reasoning from the perspective of information flow and propose a consise and organized perception approach, drawing inspiration from the perspective of human problem-solving.
@@
Benefiting from LLMs’ strong logical reasoning ability, some methods [10, 13, 44] seek to encourage LLMs to generate reasoning steps explicitly and then produce results in a single stage, while some other methods seek to perform inference at multiple times to complete the tasks [45, 24]. Several recent works, such as LOGIC-LM [30], integrate LLMs with symbolic reasoning to improve logical problem-solving. LOGIC-LM first utilizes LLMs to translate a natural language problem into a symbolic formulation. Afterward, a deterministic symbolic solver performs inference on the formulated problem. Selection-Inference [11] alternates between selection and inference to generate a series of casual reasoning steps, and LAMBADA [25] develops a backward chaining algorithm to decompose reasoning into sub-modules. In addition to prompting methods, some works aim to fine-tune LLMs to produce the final answer directly, keeping reasoning implicit [9, 27]. In contrast to their works from the perspective of how to plan that encourage or teach LLMs how to solve complex logical problems, we introduce the effective and compelling "Concise" and "Organized" strategies to reduce the difficulty of LLMs’ reasoning planning and better elicit their reasoning abilities, which can be considered as an alternative angle: reducing the difficulty of planning, or in other words, easy to plan.
@@ 3 Saliency score analysis
As a prevalent paradigm for interpretation, the concept of information flow can be instrumental in dynamically identifying critical interactions amongst tokens [12, 19, 40]. In this section, we leverage the information flow analysis methodology, predicated upon saliency scores derived from [40], to delve deeper into the reasons behind the pronounced degradation in LLMs’ performance when confronted with disordered or irrelevant information. We perform analysis on the ProofWriter [35] dataset based on Llama-2-13B-Chat [38], and the detailed saliency score definition and analysis are listed in Appendix A.1. Figure 2(a) shows reasoning steps and saliency score analysis on a concise and organized example, which is consistent with the example in Figure 1. For comparison, the premises of example in Figure 2(b) are shuffled for analysis the impact of disordered information. Multiple irrelevant premises are added in the example in Figure 2(c), but the order of the relevant premises is consistent with the example in Figure 2(a), for analysis the impact of irrelevant information. From Figure 2, we can observe three phenomena. (i) Model can identify the correct entry point of the reasoning path during the initial step when confronted with concise and organized reasoning content, which is highlighted by the green box in Figure 2(a). In contrast, when the model encounters input that is presented in a disordered sequence or contains irrelevant information, it becomes markedly challenging for the model to ascertain an appropriate entry point for reasoning at the initial step, which are highlighted by the red boxes in Figure 2(b)(c). (ii) The information flow from the previous step to the current step is salient, as clearly depicted by the diagonal lines in Figure 2(a)(b)(c). This preference for the previous step excessively focuses attention there, complicating the allocation of sufficient effort to identify the subsequent correct step. Furthermore, it may lead to the generation of non-existent premises to cater to the content of the previous step, especially when confronted with disordered and irrelevant content. In Figure 2(b), "if sb is nice then they are not young" is a fake generated premise, which can be considered as hallucination. (iii) Finally, Figure 2(c) identifies an adverse effect of irrelevant information on reasoning, with the area highlighted by purple box clearly showing a pronounced manifestation of this impact. This leads to model becoming distractible, causing it to incorporate irrelevant content into the reasoning path, and ultimately resulting in reasoning failure.
Figure 2: (a)(b)(c) Saliency score analysis on an example of ProofWriter based on shallow layers of Llama-2-13B-Chat. The horizontal coordinate contains the outputs the model generates step by step, and the vertical coordinate contains the inputs and outputs. Values in the plot represent saliency scores from column to row, normalized by each column. (a) A concise and organized example as base. (b) An example with disordered information for comparison. (c) An example with irrelevant information for comparison. (d) Saliency score analysis on ProofWriter based on shallow and deep layers of Llama-2-13B-Chat. "Base", "Disordered", and "Irrelevant" respectively denote the samples corresponding to the three scenarios depicted in (a)(b)(c). (d)(1) The saliency scores from the ground-truth reasoning entrance to the first reasoning step. (d)(2) The proportion of samples with the highest saliency score from the ground-truth reasoning entrance to the first reasoning step. (d)(3) The saliency scores from the previous two steps to the current step. (d)(4) The proportion of information flow from relevant and irrelevant information when contains irrelevant information.
@@
Further, we conduct information flow analysis on 1200 samples in the ProofWriter dataset to understand why LLMs’ performance significantly decreases when faced with disordered or irrelevant information, from a more holistic viewpoint. Drawing on  [40], we conduct a comprehensive analysis of information flow across both the shallow and deep layers. As shown in Figure 2(d)(1), when inputs are concise and organized, the saliency score from the ground-truth reasoning entrance to the first reasoning step is significantly higher than when faced with disordered or irrelevant content. Figure 2(d)(2) shows that the proportion of samples with the highest saliency score from the ground-truth reasoning entrance to the first reasoning step is still the highest when inputs are concise and organized. Intuitively, such a preference for the reasoning entry aligns with the problem-solving processes observed in humans because locating information within organized and concise sets is considerably more straightforward than navigating through a pile of complex information. Figure 2(d)(3) shows the information flow from the previous two steps to the current step, which is also salient as we observed in the diagonal lines in Figure 2(a)(b)(c). In additional, irrelevant information has dispersed the flow of information as illustrated in Figure 2(d)(4), which is very similar to the inertia of humans. However, in contrast to LLMs, the disordered and irrelevant content does not significantly decrease human performance, as humans have a propensity to distill the most relevant information and systematically organize their thoughts, aiding them in responding to questions. Thus, draw inspiration from the aspect of human problem-solving, we introduce two effective and compelling strategies including "Concise" and "Organized" perception to better elicit LLMs’ reasoning abilities.

### v5

@@ Abstract
Exploiting large language models (LLMs) to tackle reasoning has garnered growing attention. It still remains highly challenging to achieve satisfactory results in complex logical problems, characterized by plenty of premises within the context and requiring multi-hop reasoning. In particular, the reasoning capabilities of LLMs are brittle to disorder and distractibility. In this work, we first examine the mechanism from the perspective of information flow and reveal that LLMs confront difficulties akin to human-like cognitive biases when dealing with disordered and irrelevant content in reasoning tasks. However, in contrast to LLMs, disordered and irrelevant content does not significantly decrease human performance, as humans have a propensity to distill the most relevant information and systematically organize their thoughts, aiding them in responding to questions. Stem from that, we further propose a novel reasoning approach named Concise and Organized Perception (COP). COP carefully analyzes the given statements to identify the most pertinent information while eliminating redundancy efficiently. It then prompts the LLMs in a more organized form that adapts to the model’s inference process. By perceiving concise and organized context, the reasoning abilities of LLMs can be better elicited. Extensive experimental results on several popular logical benchmarks (ProofWriter, PrOntoQA, PrOntoQA-OOD, and FOLIO) and mathematical benchmark (DI-GSM) show that COP significantly outperforms previous state-of-the-art methods.
Concise and Organized Perception Facilitates Reasoning in Large Language Models
Junjie Liu1, Shaotian Yan1, Chen Shen1, Zhengdong Xiao1, Liang Xie2,1, Wenxiao Wang3,1, Jieping Ye1 1Alibaba Cloud Computing 2Zhejiang University of Technology 3Zhejiang University Corresponding author: Chen Shen (zjushenchen@gmail.com)
@@ 1 Introduction
The field of large language models (LLMs) has witnessed significant progress in complex reasoning with the advent of Chain-of-thought (CoT) prompting (Wei et al., 2022) and a series of related works (Qiao et al., 2023; Kojima et al., 2022; Zhou et al., 2023; Yao et al., 2024; Besta et al., 2024). These breakthroughs have yielded remarkable achievements in various applications, including arithmetic, commonsense, symbolic reasoning, etc. (Talmor et al., 2019; Austin et al., 2021; Chen et al., 2021; Hendrycks et al., 2021; Gao et al., 2023; Bubeck et al., 2023), and have sparked widespread enthusiasm within the community to continuously explore the immense potential of LLMs in tackling complex reasoning tasks. However, the performance of LLMs drastically decreases when handling intricate tasks characterized by plenty of premises within the prompt and requiring multi-hop reasoning. One primary issue is distractibility (Shi et al., 2023), where the reasoning capabilities of LLMs are highly susceptible to deterioration when confronted with irrelevant context. Another failure mode that has garnered significant attention recently is disorder (Saparov and He, 2023; Chen et al., 2024), where the performance of LLMs Severely drops when the premise order does not align with the context required in intermediate reasoning steps. Figure 1(a) shows an example from ProofWriter (Tafjord et al., 2021), when the given premises are disordered (in deductive reasoning, disorder can be further understood as that the premises are not in the same order as the ground truth proof) and contain much information irrelevant to the reasoning, the models face a higher risk of selecting the wrong information at some stage. This often leads to an incomplete proof and subsequently an incorrect answer. Figure 1(c) illustrates such a misleading step, where the original CoT selects the wrong reasoning path (highlighted in red). This observation indicates that LLMs usually struggle with proof planning when irrelevant and disordered content hinders, as also revealed in some contemporaneous works(Berglund et al., 2023; Shi et al., 2023; Chen et al., 2024).
@@
Figure 1: (a) A 5-hop example of ProofWriter dataset, showcasing plenty of premises.Some premises are omitted for brevity. (b) Corresponding reconstruction of concise and organized perception. Superscript serial numbers represent the logical orders according to the gold proof. The concise input contains only relevant information but lacks organizational structure. In contrast, the organized input arranges statements consistently with the gold reasoning path, albeit including redundant information. (c) LLMs outputs. (d) Results of a confirmatory experiment.
@@
In this work, we first briefly investigate the underlying mechanism of the influence of disordered and irrelevant content on reasoning. Specifically, we adopt neuron saliency score analysis, which is an important approach for pinpoint the information flow and the crucial interactions between tokens (Dai et al., 2022; Hao et al., 2021; Wang et al., 2023), and we conclude three phenomena. (i) LLMs struggle to identify the correct entry point of the reasoning path when faced with disorder and distraction. (ii) The current step always highly focuses on the previous step, to the extent that it might even make up non existing premises to accommodate the preceding step. Consequently, it is challenging in allocating sufficient energy to identify the most accurate step. (iii) In addition, salient information flow from irrelevant information renders models prone to distractions, inadvertently causing them to focus on irrelevant content, which ultimately leads to failures in reasoning.
@@
The information flow analysis reveals how irrelevant and disorganized information can affect model reasoning, and suggests the inertia of LLMs when tackling complex tasks is very similar to the one during human problem-solving process (Hagendorff et al., 2023; Jones and Steinhardt, 2022). However, in contrast to LLMs, the disordered and irrelevant content does not significantly decline human performance, which benefits from the fact that humans tend to distill the most relevant information and organize their thoughts in an orderly manner, such as constructing a mind map, in advance. This allows them to address the question accurately by referring to the mind map (Girotto et al., 1997; Dekeyser et al., 2000).
@@
Arise from that, we propose a novel reasoning approach named Concise and Organized Perception (COP). Specifically, COP initially performs capturing of locally-related premise segments among the given premises, with the intent to facilitate an initial comprehension of the input context. Next, COP leverages the query question as an anchor to integrate relevant pieces of locally-related premises generated by the first step, creating a tree-like mind map structure that presents global information in an orderly manner and can eliminate irrelevant information. Subsequently, LLMs are prompted by the reconstructed context, which are organized in a progressively ordered manner from the mind map to better adapt to the inference process of the model.
@@
We believe such reconstruction perceives more concise and organized information, which noticeably reduces the difficulty of LLMs’ planning and better elicits the reasoning ability. Figure 1(b)(c) shows an example where LLMs are empowered to obtain the correct answer. We further conducted a simple confirmatory experiment by randomly selecting 196 samples and reconstructing the context based on the provided ground-truth proofs, as shown in Figure 1(b)  [^1] This implementation is merely demonstrative, and differs from the actual method as no ground-truth can be utilized.. The results in Figure 1(d) demonstrate that combining our approach with the CoT baseline yields a 100% relative improvement (35.9% vs 71.9%) in a 5-hop setting. The results also indicate the complementarity between concise and organized perception.
We conduct extensive experiments on several popular logical benchmarks (ProofWriter, PrOntoQA, PrOntoQA-OOD), real-world complex logical benchmark (FOLIO) and mathematical benchmark (DI-GSM). The experimental results demonstrate that COP significantly outperforms previous state-of-the-art methods. Specially, COP surpasses the CoT baseline by 9% on the FOLIO benchmark.
@@ 2 Related work
LLMs have demonstrated impressive few-shot learning capabilities (Brown et al., 2020; Raffel et al., 2020; Chung et al., 2022; Ouyang et al., 2022; Touvron et al., 2023a). Recent work has shown that LLMs, combined with in-context learning (ICL) and chain-of-thought (CoT) prompting, are capable of reasoning to an extent (Huang and Chang, 2023; Qiao et al., 2023; Nye et al., 2021; Wei et al., 2022; Kojima et al., 2022; Lewkowycz et al., 2022). However, it still remains highly challenging to achieve satisfactory results in complex logical problems.  (Xu et al., 2023) explores the logical flaws of LLMs on logical reasoning datasets from four dimensions including answer correctness, explanation correctness, explanation completeness and explanation redundancy.  (Wan et al., 2024) proposes an automatic approach to evaluate the logical reasoning abilities of LLMs based on propositional and predicate logic, which systematically identifies poor logical rules for LLMs’ reasoning. In particular, the reasoning capabilities of LLMs are susceptible to deterioration when confronted with inputs that are either arranged in a disordered manner (Saparov and He, 2023; Chen et al., 2024) or peppered with irrelevant information (Shi et al., 2023). Specifically,  (Saparov and He, 2023) investigated how reasoning ability is affected by the traversal direction of the ontology, and  (Chen et al., 2024) found that the premise order significantly affects LLMs’ reasoning performance. In another study,  (Shi et al., 2023) observed that the performance of language models tends to decrease when irrelevant context is included in the problem statement. These phenomena aligns with the human preferences for solving logical problems (Girotto et al., 1997; Dekeyser et al., 2000; Johnson-Laird, 2001). Differing from their works, we further investigate the impact of disordered and irrelevant content on reasoning from the perspective of information flow and propose a concise and organized perception approach, drawing inspiration from the perspective of human problem-solving.
Figure 2: (a)(b)(c) Saliency score analysis on an example of ProofWriter based on shallow layers of Llama-2-13B-Chat. The horizontal coordinate contains the step by step outputs, and the vertical coordinate contains the inputs and outputs. Values in the plot represent saliency scores from column to row, normalized by each column.
Figure 3: Saliency score analysis on ProofWriter based on shallow and deep layers of Llama-2-13B-Chat. "Base", "Disordered", and "Irrelevant" respectively denote the samples corresponding to the three scenarios depicted in Figure 2 (a)(b)(c).
@@
Benefiting from LLMs’ strong logical reasoning ability, some methods (Cobbe et al., 2021; Dalvi et al., 2021; Zelikman et al., 2022) seek to encourage LLMs to generate reasoning steps explicitly and then produce results in a single stage, while some other methods seek to perform inference at multiple times to complete the tasks (Zhou et al., 2023; Jung et al., 2022). Several recent works, such as LOGIC-LM (Pan et al., 2023), integrate LLMs with symbolic reasoning to improve logical problem-solving. Selection-Inference (Creswell et al., 2023) alternates between selection and inference to generate a series of casual reasoning steps, and LAMBADA (Kazemi et al., 2023) develops a backward chaining algorithm to decompose reasoning into sub-modules. In addition to prompting methods, some works aim to fine-tune LLMs to produce the final answer directly, keeping reasoning implicit (Clark et al., 2021; Lewkowycz et al., 2022). In contrast to their works from the perspective of how to plan that encourage or teach LLMs how to solve complex logical problems, we introduce the effective and compelling "Concise" and "Organized" strategies to reduce the difficulty of LLMs’ reasoning planning and better elicit their reasoning abilities, which can be considered as an alternative angle: reducing the difficulty of planning, or in other words, easy to plan. The detailed analysis of novelty and effectiveness of COP can be found in Appendix A.3.
@@ 3 Saliency score analysis
As a prevalent paradigm for interpretation, the concept of information flow can be instrumental in dynamically identifying critical interactions amongst tokens (Dai et al., 2022; Hao et al., 2021; Wang et al., 2023). In this section, we leverage the information flow analysis methodology, predicated upon saliency scores derived from (Wang et al., 2023), to delve deeper into the reasons behind the pronounced degradation in LLMs’ performance when confronted with disordered or irrelevant information. We perform analysis on the ProofWriter (Tafjord et al., 2021) dataset based on Llama-2-13B-Chat (Touvron et al., 2023b), and the detailed saliency score definition and analysis are listed in Appendix A.4. Figure 2(a) shows reasoning steps and saliency score analysis on a concise and organized example, which is consistent with the example in Figure 1. For comparison, the premises of example in Figure 2(b) are shuffled for analysis the impact of disordered information. Multiple irrelevant premises are added in the example in Figure 2(c), but the order of the relevant premises is consistent with the example in Figure 2(a), for analysis the impact of irrelevant information. From Figure 2, we can observe three phenomena. (i) Model can identify the correct entry point of the reasoning path during the initial step when confronted with concise and organized reasoning content, which is highlighted by the green box in Figure 2(a). In contrast, when the model encounters input that is presented in a disordered sequence or contains irrelevant information, it becomes markedly challenging for the model to ascertain an appropriate entry point for reasoning at the initial step, which are highlighted by the red boxes in Figure 2(b)(c). (ii) The information flow from the previous step to the current step is salient, as clearly depicted by the diagonal lines in Figure 2(a)(b)(c). This preference for the previous step excessively focuses attention there, complicating the allocation of sufficient effort to identify the subsequent correct step. Furthermore, it may lead to the generation of non-existent premises to cater to the content of the previous step, especially when confronted with disordered and irrelevant content. In Figure 2(b), "if sb is nice then they are not young" is a fake generated premise, which can be considered as hallucination. (iii) Finally, Figure 2(c) identifies an adverse effect of irrelevant information on reasoning, with the area highlighted by purple box clearly showing a pronounced manifestation of this impact. This leads to model becoming distractible, causing it to incorporate irrelevant content into the reasoning path, and ultimately resulting in reasoning failure.
@@
Further, we conduct quantitative analysis on 1200 samples in the ProofWriter dataset to understand why LLMs’ performance significantly decreases when faced with disordered or irrelevant information, from a more holistic viewpoint. Drawing on  (Wang et al., 2023), we conduct a comprehensive analysis of information flow across both the shallow and deep layers. As shown in Figure 2(d)(1), when inputs are concise and organized, the saliency score from the ground-truth reasoning entrance to the first reasoning step is significantly higher than when faced with disordered or irrelevant content. Figure 2(d)(2) shows that the proportion of samples with the highest saliency score from the ground-truth reasoning entrance to the first reasoning step is still the highest when inputs are concise and organized. Intuitively, such a preference for the reasoning entry aligns with the problem-solving processes observed in humans because locating information within organized and concise sets is considerably more straightforward than navigating through a pile of complex information. Figure 2(d)(3) shows the information flow from the previous two steps to the current step, which is also salient as we observed in the diagonal lines in Figure 2(a)(b)(c). In additional, irrelevant information has dispersed the flow of information as illustrated in Figure 2(d)(4), which is very similar to the inertia of humans. However, in contrast to LLMs, the disordered and irrelevant content does not significantly decrease human performance, as humans have a propensity to distill the most relevant information and systematically organize their thoughts, aiding them in responding to questions. Thus, draw inspiration from the aspect of human problem-solving, we introduce two effective and compelling strategies ("Concise" and "Organized") to better elicit LLMs’ reasoning abilities.